The molecular weight of an oxide of an element is 44. The equivalent weight of the element is 14. The atomic weight of the element is

Grade:11

1 Answers

Arun
5 years ago
Valency = Approximate atomic mass/Equivalent weight of the metal

Valency = 44/14= 44/14

Atomic weight = Equivalent weight of the metal X Valency

Atomic weight = 14 x 44/14= 44